From 20 tickets marked with the first 20 numerals, one is drawn at random. What is the chance that it will be a multiple of 3 or of 7?
Answer by edjones(8007)

There are 6 multiples of three and 2 multiples of seven.
6+2=8
8/20=2/5 the chance that it will be a multiple of 3 or of 7.
.
Ed